# Category-free Out-of-Distribution Node Detection with Feature Resonance

### v1

@@ 5 Related Works
OOD detection methods are generally categorized into entropy-based, density-based, and representation-based approaches. Entropy-based methods such as Maximum Softmax Probability (MSP) (Hendrycks & Gimpel, 2016), Energy (Liu et al., 2020), and other methods (Liang et al., 2017; Bendale & Boult, 2016; Hendrycks et al., 2018; Geifman & El-Yaniv, 2019; Malinin & Gales, 2018; Jeong & Kim, 2020; Chen et al., 2021; Wei et al., 2021; Ming et al., 2022b, a) compute OOD scores from class distributions. Still, they rely heavily on multi-category labels, which limits their use in category-free settings. Density-based methods, such as Mahalanobis distance (Lee et al., 2018) and residual flow models (Zisselman & Tamar, 2020), estimate sample probabilities based on their distribution but struggle with handling high-dimensional data and complex relationships (Ren et al., 2019; Serrà et al., 2019). Representation-based methods, including KNN (Sun et al., 2022) and NNGuide (Park et al., 2023), focus on differentiating OOD and ID nodes by analyzing learned embeddings in feature space. However, they still need a pre-trained multi-category ID classifier. In contrast, SSD (Sehwag et al., 2021) is an outlier detector that leverages self-supervised representation learning and Mahalanobis distance-based detection on unlabeled ID data.
Category-free OOD Detection in Graphs.
Category-free OOD detection in graphs aims to identify OOD nodes without relying on multi-category labels, posing unique challenges for traditional methods. Entropy-based methods, such as OODGAT (Song & Wang, 2022) and GNNSafe (Wu et al., 2023), depend on classifier outputs and are not suitable for category-free settings. Representation-based methods, including EnergyDef (Gong & Sun, 2024), aim to generate synthetic OOD nodes but often fail to capture the true features of real OOD nodes. Graph anomaly detection methods, like DOMINANT (Ding et al., 2019) and SL-GAD (Zheng et al., 2021), detect general anomalies through reconstruction errors, but they struggle to distinguish between OOD nodes and general anomalies. Recent works such as (Li et al., 2022; Bazhenov et al., 2022; Liu et al., 2023; Ding & Shi, 2023) explore graph-level OOD detection but can not be directly applied to node-level OOD detection due to the complexity of node dependencies.
@@ 6 Conclusion
In this paper, we introduce the concept of Feature Resonance for category-free OOD detection, demonstrating that unknown ID samples undergo more substantial representation changes compared to OOD samples during the optimization of known ID samples, even in the absence of multi-category labels. To effectively capture this phenomenon, we propose a label-independent, micro-level proxy that measures feature vector movements in a single training step. Building on this, we present the RSL framework, which integrates the micro-level feature resonance with synthetic OOD node generation via SGLD, enhancing OOD detection performance and offering an efficient and practical solution for category-free OOD node detection.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1 Notations
$\boldsymbol{X}$, $\boldsymbol{Y}$ the input space and the label space.
@@
$\mathbb{P}_{\boldsymbol{XY}}$ the joint data distribution for ID data..
@@
^$\widehat{\nabla},\tau$ the average gradients on labeled ID data, uncertainty score.
@@
$\boldsymbol{X}_{\text{wild}}^{\text{in}},\boldsymbol{X}_{\text{wild}}^{\text{out}}$ inliers and outliers in the wild dataset.
$\boldsymbol{X}^{\text{in}},\boldsymbol{X}_{\text{wild}}$ labeled ID data and unlabeled wild data.
$n,m$ size of $\boldsymbol{X}^{\text{in}}$, size of $\boldsymbol{X}_{\text{wild}}$
@@
$\boldsymbol{X}_{T}$ wild data whose uncertainty score higher than threshold $T$
@@
$R_{\boldsymbol{X}}(h_{\theta_{t}})$ the empirical risk w.r.t. predictor $h_{\theta_{t}}$ over data $\boldsymbol{X}$
$R_{\mathbb{P}_{\boldsymbol{XY}}}(h_{\theta_{t}})$ the risk w.r.t. predictor $h_{\theta_{t}}$ over distribution $\mathbb{P}_{\boldsymbol{XY}}$.
@@ A.2 Definitions
Definition A.1.
($\beta$smooth).We say a loss function $\ell(h_{\theta_{t}}(\mathbf{x}),y)$ (defined over $\boldsymbol{X}\times\boldsymbol{Y})$ is $\beta$smooth, if forany $\mathbf{x}\in{\boldsymbol{X}}$ and $y\in\boldsymbol{Y}$
@@
Definition A.3.
$(\gamma,\xi)$discrepancy). We say a wild distribution $\mathbb{P}_{wild}$ has $(\gamma,\xi)$discrepancy w.r.t. an ID joint distribution $\mathbb{P}_{in}$ n, if $\gamma>\min_{\mathbf{w}\in\mathcal{W}}$ $R_{\mathbb{P}_{XY}}(h_{\theta})$ and for any parameter $\mathbf{w}\in\mathcal{W}$ satisfying that $R_{\mathbb{P},\boldsymbol{X}\boldsymbol{Y}}(h_{\theta_{t}})\leq\gamma$ should meet the following condition
@@ A.3 Assumptions
  • $\ell(h_{\theta_{t}}(\mathbf{x}),y)\geq 0$ and $\ell(h_{\theta_{t}}(\mathbf{x}),y)$ is $\beta_{1}$smooth;
  • $\sup_{(\mathbf{x},y)\in\boldsymbol{X}\times\boldsymbol{Y}}\|\nabla\ell(h_{\theta_{0}}(\mathbf{x}),y)\|_{2}=b_{1};$
  • $\sup_{(\mathbf{x},y)\in\boldsymbol{X}\times\boldsymbol{Y}}\ell(h_{\theta_{0}}(\mathbf{x}),y)=B_{1}.$
Assumption 2. $\ell(\mathbf{f}(\mathbf{x}),\widehat{y}_{\mathbf{x}})\leq\min_{y\in\boldsymbol{Y}}\ell(\mathbf{f}(\mathbf{x}),y)$ , where $\widehat{y}_{\mathbf{x}}$ returns the closest vector of the predictor $\mathbf{f}$’s output on $\mathbf{x}$
@@ A.4 Constants in Theory
$M^{\prime}=2(\beta_{1}r_{1}+b_{1})^{2}$ the upper bound of gradient-based filtering score (Du et al., 2024a)
@@
$\delta(T,t)$ the main error in 8
@@ Appendix B Main Theorems
Theorem B.1.
If Assumptions 1 and 2 hold, $\mathbb{P}_{wild}$ has $(\gamma,\xi)$discrepancy w.r.t. $\mathbb{P}_{xy}$ ,and there exists $\eta\in(0,1)$ s.t. $\Delta=(1-\eta)^{2}\xi^{2}-8\beta_{1}R_{in}^{*}>0$, then for
@@ C.1 Proof of Theorem 1
Step 1. With the probability at least $01-\frac{7}{3}\delta>$ 0
@@
This can be proven by Lemma 7 in (Du et al., 2024a) and following inequality
$\mathbb{E}_{\tilde{\mathbf{x}}_{i}\sim\mathcal{S}_{wild}^{in}}\tau_{i}\geq\mathbb{E}_{\tilde{\mathbf{x}}_{i}\sim\boldsymbol{X}_{wild}^{m}}\left\|\nabla\ell(h_{\theta_{\boldsymbol{X}^{m}}}(\tilde{\mathbf{x}}_{i}),\widehat{{h}}_{\theta_{\boldsymbol{X}^{m}}}(\tilde{\mathbf{x}}_{i}))-\mathbb{E}_{(\mathbf{x}_{j},y_{j})\sim\boldsymbol{X}^{m}}\nabla\ell(h_{\theta_{\boldsymbol{X}^{m}}}(\mathbf{x}_{j}),y_{j})\right\|_{2}^{2},$
@@
$\mathbb{E}_{\tilde{\mathbf{x}}_{i}\sim\boldsymbol{X}_{\mathrm{wild}}}\tau_{i}=\frac{|\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{in}}|}{|\boldsymbol{X}_{\mathrm{wild}}|}\mathbb{E}_{\tilde{\mathbf{x}}_{i}\sim\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{in}}}\tau_{i}+\frac{|\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{out}}|}{|\boldsymbol{X}_{\mathrm{wild}}|}\mathbb{E}_{\tilde{\mathbf{x}}_{i}\sim\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{out}}}\tau_{i}.$
@@
Under the condition in Theorem 5 in (Du et al., 2024a), with the probability at least $\frac{97}{100}-\frac{7}{3}\delta>0$
$\mathbb{E}_{\tilde{\mathbf{x}}_{i}\sim\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{out}}\tau_{i}}$ $\leq\frac{m}{|\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{out}}|}\big{[}\frac{98\eta^{2}\xi^{2}}{100}-\frac{|\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{in}}|}{m}8\beta_{1}R_{\mathrm{in}}^{*}-\frac{|\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{in}}|}{m}\epsilon(n,m)\big{]}$
$\leq\frac{m}{|\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{out}}|}\big{[}\frac{98\eta^{2}\xi^{2}}{100}-8\beta_{1}R_{\mathrm{in}}^{*}-\epsilon(n,m)\big{]}$
@@
Step 4. Under the conditions in Theorem 5 in (Du et al., 2024a) and Proposition D.4, with the probability at least $\frac{97}{100}-\frac{7}{3}\delta>0$
|$\frac{|\{\tilde{\mathbf{x}}_{i}\in\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{out}}:\tau_{i}\leq T\}|}{|\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{out}}|}\leq\frac{1-\min\{1,\Delta(n,m)\}}{1-T/(\frac{\sqrt{2}}{2t\alpha-1})^{2}},$ (13)
@@
|$\frac{|\{\tilde{\mathbf{x}}_{i}\in\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{out}}:\tau_{i}<T\}|}{|\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{out}}|}=P(Z(I)<T/(\frac{\sqrt{2}}{2t\alpha-1})^{2})\geq\frac{\Delta(n,m)-T/(\frac{\sqrt{2}}{2t\alpha-1})^{2}}{1-T/(\frac{\sqrt{2}}{2t\alpha-1})^{2}}.$ (14)
@@
|$\frac{|\{\tilde{\mathbf{x}}_{i}\in\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{out}}:\tau_{i}\leq T\}|}{|\boldsymbol{X}_{\mathrm{wild}}^{\mathrm{out}}|}\leq\frac{\epsilon+(\frac{\sqrt{2}}{2t\alpha-1})^{2}\epsilon^{\prime}(n,m)}{(\frac{\sqrt{2}}{2t\alpha-1})^{2}-T},$ (15)
@@ C.2 Proof of Theorem 2
Appendix D Necessary Propositions.
@@ D.1 Boundedness
$\begin{gathered}\operatorname*{sup}_{\mathbf{w}\in\mathcal{W}}\operatorname*{sup}_{(\mathbf{x},y)\in\boldsymbol{X}\times\boldsymbol{Y}}\|\nabla\ell(h_{\theta_{t}}(\mathbf{x}),y)\|_{2}\leq\beta_{1}r_{1}+b_{1}={\sqrt{M^{\prime}/2}},\\ \sup_{\mathbf{w}\in\mathcal{W}}\sup_{(\mathbf{x},y)\in\boldsymbol{X}\times\boldsymbol{Y}}\ell(h_{\theta_{t}}(\mathbf{x}),y)\leq\beta_{1}r_{1}^{2}+b_{1}r_{1}+B_{1}=M,\\ \end{gathered}$
@@
Proposition D.3.
If Assumption 1 holds, for any labeled data $\boldsymbol{X}$ and distribution $\mathbb{P}$.
$\left\|\nabla R_{\boldsymbol{X}}(h_{\theta_{t}})\right\|_{2}^{2}\leq 2\beta_{1}R_{\boldsymbol{X}}(h_{\theta_{t}}),\quad\forall\mathbf{w}\in\mathcal{W},$ (16)
@@
Proof. Jensen’s inequality implies that $R_{S}(h_{\theta_{t}})$ and $R_{\mathbb{P}}(\mathbf{f}_{\mathrm{w}})$ are $\beta_{1}$smooth.Then Proposition 2 implies the results.
Proposition D.4.
@@
Appendix E A Straightforward Explanation of Feature Resonance
$\begin{split}&\Delta h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})\\ &=h_{\theta_{t+1}}(\tilde{\mathbf{x}}_{i})-h_{\theta_{t}}(\tilde{\mathbf{x}}_{i})\\ &=-\alpha\ {\tilde{\mathbf{x}}_{i}\ \nabla_{\theta_{t}}\ell({\boldsymbol{X}}_{\text{known}})}\\ &=2\alpha\mathbb{E}(\underbrace{\tilde{\mathbf{x}}_{i}{\boldsymbol{X}}_{\text{known}}^{\top}}_{\text{Term 1}}(\underbrace{({\boldsymbol{X}}_{\text{known}}\mathbf{W}^{\top}_{t})-\mathbf{1}^{\top}e)}_{\text{Term 2}})\end{split}$ (18)
where $\alpha$ is the learning rate. Term 1 in the Equation 18 illustrates that when the features of $\tilde{\mathbf{x}}_{i}$ are consistent with the overall features of the labeled ID nodes ${\boldsymbol{X}}_{\text{known}}$, the representation of $\tilde{\mathbf{x}}_{i}$ undergoes a more significant change. Meanwhile, since term 2 in the Equation 18 and $\tilde{\mathbf{x}}_{i}$ are independent, the choice of the target vector can be arbitrary. It is highly suitable for category-free OOD detection scenarios, requiring no multi-category labels as ground truth.
Appendix F Experiment Details

### v2

@@ 4 Related Work
General OOD Detection Methods. OOD detection methods fall into three main categories: entropy-based, density-based, and representation-based approaches. Entropy-based methods (e.g., MSP (Hendrycks and Gimpel, 2016), Energy (Liu et al., 2020), and others (Liang et al., 2017; Bendale and Boult, 2016; Hendrycks et al., 2018; Geifman and El-Yaniv, 2019; Malinin and Gales, 2018; Jeong and Kim, 2020; Chen et al., 2021; Wei et al., 2021; Ming et al., 2022b, a)) compute scores from class distributions but rely heavily on labeled data, making them less suitable for label-free scenarios. Density-based methods (Lee et al., 2018; Zisselman and Tamar, 2020) estimate sample likelihoods but struggle with high-dimensional, complex data (Ren et al., 2019; Serrà et al., 2019). Representation-based methods like KNN (Sun et al., 2022) and NNGuide (Park et al., 2023) operate in embedding space but still require pre-trained ID classifiers. In contrast, SSD (Sehwag et al., 2021) avoids label dependence by using self-supervised learning on unlabeled ID data.
General OOD Node Detection Methods. Entropy-based methods, such as MSP Hendrycks and Gimpel (2016), ODIN Liang et al. (2017), OE Hendrycks et al. (2018), Energy & Energy FineTune Liu et al. (2020), OODGAT (Song and Wang, 2022), GNNSafe (Wu et al., 2023), NodeSafe (Yang et al., 2025), and GRASP (Ma et al., 2024), as well as recent approaches like GOLD (Wang et al., 2025), EDBD (Um et al., 2025), and DeGEM (Chen et al., 2025), all rely on the outputs of a pre-trained classifier, making them unsuitable for unsupervised settings. Graph anomaly detection methods, like DOMINANT (Ding et al., 2019) and SL-GAD (Zheng et al., 2021), detect general anomalies through reconstruction errors, but they struggle to distinguish between OOD nodes and general anomalies.
Unsupervised OOD Node Detection Methods. Unsupervised OOD node detection in graphs aims to identify OOD nodes without relying on multi-category labels and pretext classification tasks, posing unique challenges for traditional methods. Recent works (Li et al., 2022; Bazhenov et al., 2022; Liu et al., 2023; Ding and Shi, 2023) explore graph-level OOD detection but can not be directly applied to node-level OOD detection due to the complexity of node dependencies. EnergyDef (Gong and Sun, 2024) employs a synthetic OOD node strategy for unsupervised OOD node detection, and we follow up by significantly improving OOD node detection performance in the unsupervised setting.
@@ 5 Conclusion
In this paper, we introduce the concept of Feature Resonance for unsupervised OOD node detection, demonstrating that unknown ID samples undergo more substantial representation changes compared to OOD samples during the optimization of known ID samples, even in the absence of multi-class labels. To effectively capture this phenomenon, we propose a label-independent, micro-level proxy that measures feature vector movements in a single training step. Building on this, we present the RSL framework, which integrates the micro-level feature resonance with synthetic OOD node generation via SGLD, enhancing OOD detection performance and offering an efficient and practical solution for unsupervised OOD node detection.
6 Acknowledgement
This paper is mainly supported by the NSFC under Grants (No. 62402424). Haobo Wang is also supported by the Fundamental Research Funds for the Central Universities (No. 226-2025-00004) and Zhejiang Provincial Universities (No. 226-2025-00065).
@@ A.1 Notations
$\bm{X}$, $\bm{Y}$ the input space and the label space.
@@
$\mathbb{P}_{\bm{XY}}$ the joint data distribution for ID data..
@@
$\widehat{\nabla},\tau$ the average gradients on labeled ID data, uncertainty score.
@@
$\bm{X}_{\text{wild}}^{\text{in}},\bm{X}_{\text{wild}}^{\text{out}}$ inliers and outliers in the wild dataset.
$\bm{X}^{\text{in}},\bm{X}_{\text{wild}}$ labeled ID data and unlabeled wild data.
$n,m$ size of $\bm{X}^{\text{in}}$, size of $\bm{X}_{\text{wild}}$
@@
$\bm{X}_{T}$ wild data whose uncertainty score higher than threshold $T$
@@
$R_{\bm{X}}(h_{\theta_{t}})$ the empirical risk w.r.t. predictor $h_{\theta_{t}}$ over data $\bm{X}$
$R_{\mathbb{P}_{\bm{XY}}}(h_{\theta_{t}})$ the risk w.r.t. predictor $h_{\theta_{t}}$ over distribution $\mathbb{P}_{\bm{XY}}$.
@@ A.2 Definitions
Definition 3.
($\beta$ -smooth).We say a loss function $\ell(h_{\theta_{t}}(\mathbf{x}),y)$ (defined over $\bm{X}\times\bm{Y})$ is $\beta$ -smooth, if forany $\mathbf{x}\in{\bm{X}}$ and $y\in\bm{Y}$
@@
Definition 5.
$(\gamma,\xi)$ -discrepancy). We say a wild distribution $\mathbb{P}_{wild}$ has $(\gamma,\xi)$ -discrepancy w.r.t. an ID joint distribution $\mathbb{P}_{in}$ n, if $\gamma>\min_{\mathbf{w}\in\mathcal{W}}$ $R_{\mathbb{P}_{XY}}(h_{\theta})$ and for any parameter $\mathbf{w}\in\mathcal{W}$ satisfying that $R_{\mathbb{P},\bm{X}\bm{Y}}(h_{\theta_{t}})\leq\gamma$ should meet the following condition
@@ A.3 Assumptions
  • $\ell(h_{\theta_{t}}(\mathbf{x}),y)\geq 0$ and $\ell(h_{\theta_{t}}(\mathbf{x}),y)$ is $\beta_{1}$ -smooth;
  • $\sup_{(\mathbf{x},y)\in\bm{X}\times\bm{Y}}\|\nabla\ell(h_{\theta_{0}}(\mathbf{x}),y)\|_{2}=b_{1};$
  • $\sup_{(\mathbf{x},y)\in\bm{X}\times\bm{Y}}\ell(h_{\theta_{0}}(\mathbf{x}),y)=B_{1}.$
Assumption 2. $\ell(\mathbf{f}(\mathbf{x}),\widehat{y}_{\mathbf{x}})\leq\min_{y\in\bm{Y}}\ell(\mathbf{f}(\mathbf{x}),y)$ , where $\widehat{y}_{\mathbf{x}}$ returns the closest vector of the predictor $\mathbf{f}$’s output on $\mathbf{x}$
@@ A.4 Constants in Theory
$M^{\prime}=2(\beta_{1}r_{1}+b_{1})^{2}$ the upper bound of gradient-based filtering score [Du et al., 2024a]
@@
$\delta(T,t)$ the main error in Eq. 7
@@ Appendix B Main Theorems
Theorem 3.
If Assumptions 1 and 2 hold, $\mathbb{P}_{wild}$ has $(\gamma,\xi)$ -discrepancy w.r.t. $\mathbb{P}_{xy}$ ,and there exists $\eta\in(0,1)$ s.t. $\Delta=(1-\eta)^{2}\xi^{2}-8\beta_{1}R_{in}^{*}>0$, then for
@@ C.1 Proof of Theorem 1
Step 1. With the probability at least $1-\frac{7}{3}\delta>0$
@@
This can be proven by Lemma 7 in [Du et al., 2024a] and following inequality
$\mathbb{E}_{\tilde{\mathbf{x}}_{i}\sim\mathcal{S}_{wild}^{in}}\tau_{i}\geq\mathbb{E}_{\tilde{\mathbf{x}}_{i}\sim\bm{X}_{wild}^{m}}\left\|\nabla\ell(h_{\theta_{\bm{X}^{m}}}(\tilde{\mathbf{x}}_{i}),\widehat{{h}}_{\theta_{\bm{X}^{m}}}(\tilde{\mathbf{x}}_{i}))-\mathbb{E}_{(\mathbf{x}_{j},y_{j})\sim\bm{X}^{m}}\nabla\ell(h_{\theta_{\bm{X}^{m}}}(\mathbf{x}_{j}),y_{j})\right\|_{2}^{2},$
@@
$\mathbb{E}_{\tilde{\mathbf{x}}_{i}\sim\bm{X}_{\mathrm{wild}}}\tau_{i}=\frac{|\bm{X}_{\mathrm{wild}}^{\mathrm{in}}|}{|\bm{X}_{\mathrm{wild}}|}\mathbb{E}_{\tilde{\mathbf{x}}_{i}\sim\bm{X}_{\mathrm{wild}}^{\mathrm{in}}}\tau_{i}+\frac{|\bm{X}_{\mathrm{wild}}^{\mathrm{out}}|}{|\bm{X}_{\mathrm{wild}}|}\mathbb{E}_{\tilde{\mathbf{x}}_{i}\sim\bm{X}_{\mathrm{wild}}^{\mathrm{out}}}\tau_{i}.$
@@
Under the condition in Theorem 5 in [Du et al., 2024a], with the probability at least $\frac{97}{100}-\frac{7}{3}\delta>0$
$\mathbb{E}_{\tilde{\mathbf{x}}_{i}\sim\bm{X}_{\mathrm{wild}}^{\mathrm{out}}\tau_{i}}$ $\leq\frac{m}{|\bm{X}_{\mathrm{wild}}^{\mathrm{out}}|}\big[\frac{98\eta^{2}\xi^{2}}{100}-\frac{|\bm{X}_{\mathrm{wild}}^{\mathrm{in}}|}{m}8\beta_{1}R_{\mathrm{in}}^{*}-\frac{|\bm{X}_{\mathrm{wild}}^{\mathrm{in}}|}{m}\epsilon(n,m)\big]$
$\leq\frac{m}{|\bm{X}_{\mathrm{wild}}^{\mathrm{out}}|}\big[\frac{98\eta^{2}\xi^{2}}{100}-8\beta_{1}R_{\mathrm{in}}^{*}-\epsilon(n,m)\big]$
@@
Step 4. Under the conditions in Theorem 5 in [Du et al., 2024a] and Proposition 4, with the probability at least $\frac{97}{100}-\frac{7}{3}\delta>0$
$\frac{|\{\tilde{\mathbf{x}}_{i}\in\bm{X}_{\mathrm{wild}}^{\mathrm{out}}:\tau_{i}\leq T\}|}{|\bm{X}_{\mathrm{wild}}^{\mathrm{out}}|}\leq\frac{1-\min\{1,\Delta(n,m)\}}{1-T/(\frac{\sqrt{2}}{2t\alpha-1})^{2}},$ (11)
@@
$\frac{|\{\tilde{\mathbf{x}}_{i}\in\bm{X}_{\mathrm{wild}}^{\mathrm{out}}:\tau_{i}<T\}|}{|\bm{X}_{\mathrm{wild}}^{\mathrm{out}}|}=P(Z(I)<T/(\frac{\sqrt{2}}{2t\alpha-1})^{2})\geq\frac{\Delta(n,m)-T/(\frac{\sqrt{2}}{2t\alpha-1})^{2}}{1-T/(\frac{\sqrt{2}}{2t\alpha-1})^{2}}.$ (12)
@@
$\frac{|\{\tilde{\mathbf{x}}_{i}\in\bm{X}_{\mathrm{wild}}^{\mathrm{out}}:\tau_{i}\leq T\}|}{|\bm{X}_{\mathrm{wild}}^{\mathrm{out}}|}\leq\frac{\epsilon+(\frac{\sqrt{2}}{2t\alpha-1})^{2}\epsilon^{\prime}(n,m)}{(\frac{\sqrt{2}}{2t\alpha-1})^{2}-T},$ (13)
@@ C.2 Proof of Theorem 2
Appendix D Necessary Propositions
@@ D.1 Boundedness
[…]
@@
Proposition 3.
If Assumption 1 holds, for any labeled data $\bm{X}$ and distribution $\mathbb{P}$.
$\left\|\nabla R_{\bm{X}}(h_{\theta_{t}})\right\|_{2}^{2}\leq 2\beta_{1}R_{\bm{X}}(h_{\theta_{t}}),\quad\forall\mathbf{w}\in\mathcal{W},$ (14)
@@
Proof. Jensen’s inequality implies that $R_{S}(h_{\theta_{t}})$ and $R_{\mathbb{P}}(\mathbf{f}_{\mathrm{w}})$ are $\beta_{1}$ -smooth.Then Proposition 2 implies the results.
Proposition 4.
@@
Appendix E Experiment Details